OpenStudy (anonymous):

Write the standard form of 25x2 - 4y2 - 150x - 16y = -109 Find the coordinates of the center.

OpenStudy (anonymous):

We have to use here Completion Of Square Method..

OpenStudy (anonymous):

right..would you be able to please explain that?

OpenStudy (anonymous):

It will take time but yeah I do if you will co operate with proper..

OpenStudy (anonymous):

Tell me what you get if you divide the coefficient of x by 2??

OpenStudy (anonymous):

75

OpenStudy (anonymous):

Yes.... Do you know this formula: \[a^2 - 2ab + b^2 = (a-b)^2\]

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Ok.. So we are to reduce the terms in this formula so that we can use it... Now, take 25 common from: 25x^2 - 150x What you will get?

OpenStudy (anonymous):

x^2-75x

OpenStudy (anonymous):

You are 25 common.. How you got 75??

OpenStudy (anonymous):

sorry i mean x^2-6x

OpenStudy (anonymous):

yes now we have to solve for this; now divide the coefficient of x by 2..

OpenStudy (anonymous):

-3x

OpenStudy (anonymous):

really??

OpenStudy (anonymous):

What is coefficient of x??

OpenStudy (anonymous):

is it ... 25(x-3)^2-4(y+2)^2=100 or ((x-3)^2)/4 -((y+2)^2)/25 =1

OpenStudy (anonymous):

Wait I am checking.. I am just making him to solve and not yet solved by myself...

OpenStudy (anonymous):

are coordinates of the centre is 3,-2

OpenStudy (anonymous):

yes!! thank you!!!

OpenStudy (anonymous):

mention not